Required information

Use the following information for the Exercises below.

[The following information applies to the questions displayed below.]
  
Laker Company reported the following January purchases and sales data for its only product.
  

Date Activities Units Acquired at Cost Units sold at Retail
Jan. 1 Beginning inventory 205 units @ $ 13.00 = $ 2,665
Jan. 10 Sales 165 units @ $ 22.00
Jan. 20 Purchase 140 units @ $ 12.00 = 1,680
Jan. 25 Sales 145 units @ $ 22.00
Jan. 30 Purchase 310 units @ $ 11.50 = 3,565
Totals 655 units $ 7,910 310 units

rev: 09_15_2017_QC_CS-99723

Exercise 5-5A Periodic: Inventory costing LO P3

Required:

The Company uses a periodic inventory system. For specific identification, ending inventory consists of 345 units, where 310 are from the January 30 purchase, 5 are from the January 20 purchase, and 30 are from beginning inventory. Determine the cost assigned to ending inventory and to cost of goods sold using (a) specific identification, (b) weighted average, (c) FIFO, and (d) LIFO.
Answer #1

Calculate ending inventory

Specific identification Weighted average FIFO LIFO
Ending inventory (30*13+5*12+3565) = 4015 7910/655*345 = 4166 (310*11+35*12) = 3985 (205*13+140*12) = 4345